# Kinetically Consistent Coarse Graining using Kernel-based Extended Dynamic Mode Decomposition

### v1

@@ 3 Methods
3.1 Overview and Notation
Successful application of the gEDMD method requires selecting proper and expressive basis functions. This is not a trivial task and can highly depend on the dynamics of interest. To alleviate this, use of kernel methods has been proposed in Ref. 32 . The kgEDMD method is based on choosing the basis functions as features of a reproducing kernel $k(x,y)$ at the training data:
$\Phi=\left[k(\cdot,x_{1}),k(\cdot,x_{2}),...,k(\cdot,x_{m})\right]^{\top}.$ (24)
We consider $\Phi$ as a vector-valued function, that is, we write
$\Phi(x)=\left[k(x,x_{1}),k(x,x_{2}),...,k(x,x_{m})\right]^{\top}$ (25)
for the $m$-dimensional vector obtained by evaluating all the kernel features at $x$. Since kernels are two-argument functions, we can take their derivatives with respect to both arguments, and in fact we will need to distinguish these derivatives when formulating our methods. In general, we simply write $\nabla_{1}k$ and $\nabla_{2}k$ for gradients of the kernel function with respect to its first and second argument. Similarly, we write $\nabla_{1}\Phi,\,\nabla_{2}\Phi$ for the matrix-valued functions we obtain by assembling the gradients of all kernel features:
$\nabla_{1}\Phi(x)$ $=\begin{bmatrix}\nabla_{1}k(x,x_{1})\\ \vdots\\ \nabla_{1}k(x,x_{m})\end{bmatrix}\in\mathbb{R}^{m\times d},$ $\nabla_{2}\Phi(x)$ $=\begin{bmatrix}\nabla_{2}k(x,x_{1})\\ \vdots\\ \nabla_{2}k(x,x_{m})\end{bmatrix}\in\mathbb{R}^{m\times d}.$ (26)
One of the widely used kernel functions is the Gaussian kernel given by
$k(x_{i},x_{j})=\exp\left(-\frac{\|x_{i}-x_{j}\|)^{2}}{2\sigma^{2}}\right),$ (27)
where $\sigma$ is the standard deviation or bandwidth of the Gaussian kernel. If the state space is a periodic domain, we make use of a periodic Gaussian kernel instead 41
$k(x_{i},x_{j})=\exp\left(-\frac{\sum_{i=1}^{d}\sin\left(\frac{\pi}{p}(x_{i}-x_{j})\right)^{2}}{2\sigma^{2}}\right),$ (28)
where $\sigma$ is again the bandwidth of the Gaussian kernel, $p$ is the half-period of the domain, and $d$ is the dimension of system.
3.2 Kernel-based Koopman generator EDMD
Using the kernel feature basis as in Equation (24), we can seek a data-driven approximation to the eigenfunctions of the Koopman generator in the form $\psi_{i}(x)=\mathbf{u}_{i}^{\top}\Phi(x)$. As detailed in Ref. 32 , the coefficient vectors $\mathbf{u}_{i}$, are then found as solutions of the following generalized matrix eigenvalue problem
$\hat{\mathbf{A}}\mathbf{u}_{i}:=\frac{1}{2}\sum_{l=1}^{d}(\widehat{\mathbf{A}}^{(l)})^{\top}\widehat{\mathbf{A}}^{(l)}\mathbf{u}_{i}=\widehat{\lambda}_{i}\widehat{\mathbf{G}}\widehat{\mathbf{G}}\mathbf{u}_{i},$ (29)
where $d$ is the dimension of the state space. The matrices $\hat{\mathbf{A}}^{(l)}$ and $\hat{\mathbf{G}}$ are based on the reversible formulation for gEDMD, and essentially contain pairwise evaluations of the kernel and its first-order derivatives
$\begin{gathered}\left[\widehat{\mathbf{G}}\right]_{ns}=k(x_{n},x_{s})\\ \left[\widehat{\mathbf{A}}^{(l)}\right]_{ns}=\sigma_{l}(x_{n})^{\top}\nabla_{2}k(x_{n},x_{s}).\end{gathered}$ (30)
The kernelized mass matrix $\hat{\mathbf{G}}$ is often highly ill-conditioned. Therefore, we compute a reduced generator by performing a whitening transformation based on removing small eigenvalues of the mass matrix
$\widehat{\mathbf{G}}$ $=\mathbf{U}\Sigma\mathbf{U}^{T},$ $\mathbf{R}$ $=\mathbf{U}\Sigma^{-0.5}\in\mathbb{R}^{m\times r},$ $\hat{\mathbf{L}}_{r}$ $=\mathbf{R}^{T}\hat{\mathbf{A}}\hskip 2.84544pt\mathbf{R},$ (31)
in which $r\leq m$. Here, $\mathbf{R}$ is a transformation matrix mapping the original basis built by the feature map $\Phi$, to the reduced basis
$\mathbf{h}(x)=\mathbf{R}^{\top}\Phi(x)\in\mathbb{R}^{r}.$ (32)
In preparation for later, we also introduce the notation
$\nabla\mathbf{h}(x)=\left[\partial_{j}h_{l}(x)\right]_{l,j}=\mathbf{R}^{\top}\nabla_{1}\Phi(x)\in\mathbb{R}^{r\times d}$ (33)
for the gradients of all reduced features evaluated at point $x$. Diagonalization of the reduced generator matrix $\hat{\mathbf{L}}_{r}$ then provides the final estimates for the generator’s eigenvalues and eigenfunctions.
3.3 System identification
The reduced generator matrix $\hat{\mathbf{L}}_{r}$ provides an empirical model for the Koopman generator $\mathcal{L}$ on the space of whitened and rank-reduced features $\mathbf{h}$. Hence we can use this model towards other ends besides identification of the meta stable sets. Among other purposes, we can use it to identify an effective SDE governing the CG dynamics as in Equation (13), using a procedure that was already sketched in Ref. 29 . We parameterize the CG dynamics as a reversible SDE, which is justified if the original dynamics are reversible 13 . As discussed in Section 2 and in Ref. 33 , reversible systems can be characterized by the diffusion $a^{\xi}\in\mathbb{R}^{d_{\xi}\times d_{\xi}}$ and a scalar potential $F^{\xi}:\mathbb{R}^{d_{\xi}}\rightarrow\mathbb{R}$, from which the drift is then obtained by:
@@ 3.3 System identification
Thus, to identify the SDE, we need to learn these two components, which can be broken up into two separate problems. From now on, we employ a kernel $k$ on the CG space $\hat{\Omega}$, that is $k=k(z,z^{\prime})$.
3.3.1 Force matching
Knowing the full-process scalar potential field $F$ and the CG function $\xi$, the effective force can be obtained by a well-known technique called force matching 15 via the following minimization problem
$\nabla_{z}F^{\xi}=\underset{g\in L^{2}_{\nu}}{\operatorname*{arg\,min}}\frac{1}{m}\sum^{m}_{i=1}\left\|g(\xi(x_{i}))-f^{\xi}_{\text{lmf}}(x_{i})\right\|^{2},$ (35)
where $m$ is the data size and $f^{\xi}_{\text{lmf}}$ is called local mean force and defined as follows
$\begin{gathered}f^{\xi}_{\text{lmf}}=-\nabla_{x}F\cdot G^{\xi}+\nabla_{x}\cdot G^{\xi}\\ G^{\xi}=\nabla_{x}\xi[(\nabla_{x}\xi)^{T}\nabla_{x}\xi]^{-1}.\\ \end{gathered}$ (36)
In line with the kernel-based approach we follow in this work, we parameterize the optimal vector field $g$ as the gradient of a linear combination of the feature basis $\Phi$ in Equation (24)
$F^{\xi}(z)$ $=\alpha^{\top}\Phi(z),$ $g(z)$ $=\alpha^{\top}\nabla_{1}\Phi(z)=\sum_{i=1}^{m}\alpha_{i}\nabla_{1}k(z,z_{i}).$ (37)
The force matching problem thus breaks down to determination of $m$ optimal expansions coefficients $\alpha$. We stress, however, that any other parameterization could be used for this step, for example a deep learning-based method 11 .
3.3.2 Kernel-based diffusion learning
Our main contribution is a learning method for the effective diffusion, complementing the effective potential $F^{\xi}$. We first use training data of the full space dynamics and the kernel $k$ on the CG space $\hat{\Omega}$, to learn a reduced matrix representation $\hat{\mathbf{L}}_{r}$ of the full generator. Recalling Equation (15), the matrix $\hat{\mathbf{L}}_{r}$ is at the same time an approximation to the CG generator $\mathcal{L}^{\xi}$. On the CG space, the reversible generator $\mathcal{L}^{\xi}$ also satisfies the integration-by-parts formula (21),
@@ 3.3.2 Kernel-based diffusion learning
$\left\langle\mathcal{L}^{\xi}\psi,\,\varphi\right\rangle_{\nu}=-\frac{1}{2}\int_{\hat{\Omega}}\nabla_{z}\psi(z)a^{\xi}(z)\nabla_{z}\varphi(z)^{\top}\,\mathrm{d}\nu(z),$ (38)
which only features an explicit dependence on the unknown effective diffusion $a^{\xi}$, see Equation (14). Using the reduced basis $\mathbf{h}(z)=\mathbf{R}^{\top}\Phi(z)$, we parameterize the effective diffusion as
$\sigma^{\xi}_{\alpha}(z)=\sum^{r}_{l=1}\alpha_{l}h_{l}(z),$ (39)
where, in the most general case, $\alpha_{l}\in\mathbb{R}^{d_{\xi}\times d_{\xi}}$ must be matrix-valued coefficients. To arrive at a compact expression, we gather all unknown expansion coefficients $\alpha_{l}$ as a third order tensor $\alpha\in\mathbb{R}^{d_{\xi}\times d_{\xi}\times r}$, which allows us to write
$\sigma^{\xi}_{\alpha}(z)=\alpha\cdot_{|3,1}\mathbf{h}(z)=\alpha\cdot_{|3,1}\mathbf{R}^{\top}\Phi(z),$ (40)
using the notation $\mathbf{a}\cdot_{|i,j}\mathbf{b}$ to denote contraction of two tensors $\mathbf{a},\,\mathbf{b}$ over indices $i$ and $j$ of commensurate dimension. Plugging this into Equation (38) and using the data-driven estimator of Equation (30), we obtain a parametric model for the reduced generator matrix $\hat{\mathbf{L}}_{r}$
$\begin{split}\left[\hat{\mathbf{L}}^{\xi}_{\alpha}\right]&=\dfrac{1}{2m}\sum_{k=1}^{m}\nabla\mathbf{h}(z_{k})\sigma_{\alpha}^{\xi}(z_{k})\sigma_{\alpha}^{\xi}(z_{k})^{\top}\nabla\mathbf{h}^{\top}(z_{k})\\ &=\dfrac{1}{2m}\sum_{k=1}^{m}\mathbf{S}^{\alpha}(z_{k})\mathbf{S}^{\alpha}(z_{k})^{\top},\\ \mathbf{S}^{\alpha}(z)&:=\nabla\mathbf{h}(z)\sigma_{\alpha}^{\xi}(z).\end{split}$ (41)
Combining Equations (33) and (40), we can compactly write
$\mathbf{S}^{\alpha}(z)=\mathbf{R}^{\top}\nabla_{1}\Phi(z)\alpha\cdot_{|3,1}\mathbf{R}^{\top}\Phi(z),$ (42)
which is now expressed solely in terms of the kernel features $\Phi$, the whitening transformation $\mathbf{R}$, and the unknown coefficient tensor $\alpha$. We then learn the effective diffusion by matching the parameterized generator matrix $\hat{\mathbf{L}}^{\xi}_{\alpha}$ to the true reduced generator matrix $\hat{\mathbf{L}}_{r}$, therefore solving the following minimization problem:
$\alpha^{*}=\underset{\alpha\in\mathbb{R}^{d_{\xi}\times d_{\xi}\times r}}{\arg\min}\left[\left\|\hat{\mathbf{L}}_{r}-\hat{\mathbf{L}}^{\xi}_{\alpha}\right\|_{2}^{2}+\gamma\left\|\frac{\partial\sigma}{\partial z}\right\|_{2}^{2}\right],$ (43)
where $\hat{\mathbf{L}}^{\xi}_{\alpha}$ is calculated from $\alpha$ via Equations (41) and (42). To avoid over-fitting and to suppress oscillations in $\sigma$, we found that introducing the second term to regularize the derivatives of the diffusion provides the best results.
Here, the effective diffusion is represented by a full matrix. However, for simplicity, we only learn a diagonal diffusion matrix for the numerical results reported in the following section. The method can be summarized in Algorithm 1.
Algorithm 1 Learning Effective Dynamics with kgEDMD
@@
Input: full space data $\{x_{k}\}_{k=1}^{m}$ in $\mathbb{R}^{d}$, CG map $\xi$, kernel function $k$,
Truncation rule for $r$ in Equation (31), regularization $\gamma$
@@
3:     Form mass and stiffness matrices $\mathbf{G}$ and $\mathbf{A}^{(l)}$ as in Equation (30)
4:     Compute reduced generator $\hat{\mathbf{L}}_{r}$ as in Equation (31)
5:     Perform the minimization problem as in Equation (43)
6:     Form the effective diffusion $\sigma^{\xi}_{\alpha}$ as in Equation (39)
7:Force matching
8:     Compute the local mean force $f^{\xi}_{\text{lmf}}$ as in Equation (36)
9:     Perform the minimization problem as in Equation (35)
10:     Form the effective potential $F^{\xi}$ as in Equation (37)
@@ 4 Examples
To show the effectiveness of the proposed method, we apply it to a two-dimensional model system defined by the Lemon-slice potential, and to MD simulation data of the alanine dipeptide, which is a widely-used test case in molecular dynamics.
As the optimization block for diffusion learning is non-convex, we use a library function from scipy’s optimization toolbox, but solve the problem repeatedly for $50$ randomly generated initial guesses. The optimization block for the force matching, on the other hand, is done via ridge regression.
@@ 4.1.1 System introduction
As shown in previous studies, the polar angle $\phi$ is a suitable CG coordinate for this system, as it resolves all four meta stable states
@@
For this system, analytical expressions for the effective drift and diffusion along $\xi$ can be obtained by a slight modification of the results in 42 , and serve as reference values.
We use the kgEDMD algorithm on the reaction coordinate $\xi$, to identify the generator eigenvalues and meta stable states and, subsequently, to identify an effective dynamics along $\xi$ using Algorithm 1.
As a first try, we consider the Gaussian kernel for Galerkin approximation of the generator. Notice that the CG dynamics of Lemon-slice problem is periodic in the polar angle, however, a generic Gaussian kernel is not so. Thus, it cannot capture the periodicity of the system of interest. As an alternative, we also use the periodic Gaussian kernel as in Equation (28) and analyze both. The kernel bandwidth in either versions of Gaussian kernel is optimized using cross validation based on the VAMP-score 27 . Details on the VAMP-score analysis are reported in the appendix.
@@ 4.1.2 Meta-stability analysis
Figure 1 c shows the leading eigenvalues obtained from the kgEDMD model. As one notices, there are four dominant eigenvalues followed by a gap. These four eigenvalues are corresponding to the four minima in the potential field. Having determined the eigenvectors of the generator, we can perform robust Perron Cluster Cluster Analysis (PCCA+) 43 algorithm to assign to each sample point its membership to each meta stable state. Figure 1 b shows that the four potential minima are perfectly recovered in this way.
@@
A comparison of the leading eigenvalues of the reference model $\hat{\mathbf{L}}_{r}$ and the optimal matrix $\hat{\mathbf{L}}_{\alpha}^{\xi}$ is shown in Figure 1 c. Both choices of the kernel function lead to satisfactory results, the periodic kernel provides slightly higher accuracy in approximation of the generator eigenvalues. Note that the kernel bandwidth is tuned for each kernel function separately.
4.1.3 Identification of CG dynamics

### v2

@@ 3 Methods
The main proposal in this work is a method to identify a diffusion field for the coarse-grained dynamics based on data-driven approximations to the generator. We also show how the quality of the diffusion field in terms of reproducing dominant eigenvalues can be assessed, and finally, how the diffusion can be complemented to identify the complete CG dynamics. We recall that the dynamical equation in CG space is given by (12), where the drift can be written as follows because of reversibility 30
@@
3.1 Diffusion Learning
By Equation (13), the analytical effective diffusion $a^{\xi}$ is the best-approximation of the local diffusion $a^{\xi}_{\mathrm{loc}}$ by a (matrix-valued) function on the CG space. Hence, we can solve the following data-based minimization problem
$a^{\xi}=\underset{a=a(z)}{\operatorname*{arg\,min}}\frac{1}{m}\sum^{m}_{i=1}\left\|a(\xi(x_{i}))-a^{\xi}_{\mathrm{loc}}(x_{i})\right\|_{F}^{2},$ (25)
where $\|\cdot\|_{F}$ is the Frobenius norm for matrices. We parametrize the diffusion field $a^{\xi}$ element-wise as a linear combination of the reduced RFF basis
$(a^{\xi}_{\alpha})_{ij}(z)=\sum_{l=1}^{r}\alpha^{ij}_{l}h_{l}(z)=\alpha\cdot_{|3,1}\mathbf{h}(z)=\alpha\cdot_{|3,1}\mathbf{R}^{\top}\psi(z),$ (26)
where we view the coefficient array $\alpha$ as a third-order tensor of dimension $d\times d\times r$, and the symbol $\cdot_{|i,j}$ denotes contraction over indices $i$ and $j$ of two arrays. The parametrization must be symmetric, i.e. $a^{\xi}_{ij}=a^{\xi}_{ji}$, and we may also choose to set specific elements to zero, for example to enforce a diagonal diffusion field. With the parametrization (26), the minimization problem (25) becomes a linear regression problem that can be directly solved, potentially after regularization.
3.2 Recovery of Spectral Properties
After solving the minimzation problem (25), we can make use of the gEDMD method to assess the dynamical properties of the learned SDE in CG space. Using the integration-by-parts formula (19), the elements of the reduced generator matrix corresponding to the diffusion field (26) with coefficient array $\alpha$ are
@@ 3.2 Recovery of Spectral Properties
$\left\langle h_{r},\,\mathcal{L}^{\alpha}h_{s}\right\rangle_{\nu}=-\frac{1}{2}\int\nabla h_{r}(z)a^{\xi}_{\alpha}(z)\nabla h_{s}(z)^{\top}.$
In matrix notation, this leads to the following explicit formula for the parametrized generator matrix, which can be computed directly without resorting to numerical simulations of the CG dynamics
$\hat{\mathbf{L}}_{r}^{\alpha}=\sum_{i=1}^{m}\mathbf{R}^{\top}\nabla_{z}\psi(z_{i})\left[\alpha\cdot_{|3,1}\mathbf{R}^{\top}\psi(z_{i})\right]\nabla_{z}\psi(z_{i})^{\top}\mathbf{R}.$ (27)
Properties inferred from the matrix $\hat{\mathbf{L}}^{\alpha}_{r}$ can be compared to those obtained from the original gEDMD matrix $\hat{\mathbf{L}}_{r}$ estimated off the full-space simulation data. For example, diagonalization of both $\hat{\mathbf{L}}_{r}$ and $\hat{\mathbf{L}}^{\alpha}_{r}$ leads to estimates $\lambda_{i}$ and $\lambda_{i}^{\alpha}$ for the dominant generator eigenvalues, which can be systematically compared. We mainly resort to comparing dominant eigenvalues in the examples below, but we point out that a more detailed assessment is possible: for instance, by computing matrix exponentials $\exp\left(t\hat{\mathbf{L}}_{r}\right)$ and $\exp\left(t\hat{\mathbf{L}}^{\alpha}_{r}\right)$, time-correlation functions can also be evaluated.
3.3 Learning the Effective Potential
We have seen that the accuracy of the effective diffusion field largely determines the dynamical properties of the coarse grained dynamics. In order to run simulations of the CG dynamics, and to ensure thermodynamic consistency, the effective potential $F^{\xi}$ also must be learned in a parametric form. This is not the main focus of our study, hence we just point out a few options. A well-known and generally applicable technique is force matching 15 , which is based on the following minimization problem for the effective force
$\nabla_{z}F^{\xi}=\underset{g=g(z)}{\operatorname*{arg\,min}}\frac{1}{m}\sum^{m}_{i=1}\left\|g(\xi(x_{i}))-f^{\xi}_{\text{lmf}}(x_{i})\right\|^{2},$ (28)
where $f^{\xi}_{\text{lmf}}$ is called local mean force and defined as follows
$f^{\xi}_{\text{lmf}}$ $=-\nabla_{x}F\cdot G^{\xi}+\nabla_{x}\cdot G^{\xi},$ $G^{\xi}$ $=\nabla_{x}\xi[(\nabla_{x}\xi)^{\top}\nabla_{x}\xi]^{-1}.$ (29)
We point out the similarity to (25), which also led us to the name local diffusion for $a^{\xi}_{\mathrm{loc}}$. The effective potential can be parametrized as a linear combination of basis functions, such as random features, or as a deep neural network 11 . In low-dimensional CG spaces, it also possible to approximate the projected invariant distribution $\nu$ as a linear combination of kernel functions centered at the data sites, known as kernel density estimate (KDE) 36 . Since we only consider low-dimensional CG spaces here, we opt for the KDE option in the examples below.
Algorithm 1 Learning Effective Dynamics
@@ 3.3 Learning the Effective Potential
Input: full space data $\{x_{k}\}_{k=1}^{m}$ in $\mathbb{R}^{d}$, CG map $\xi$, kernel function with spectral measure $\rho$,
Truncation rule for $r$ in Equation (17), regularization parameters
@@
3:     Generate random feature basis $\psi$ as in Equation (22).
4:     Compute reduced basis $\mathbf{h}$ according to (18).
5:     Compute reduced generator matrix $\hat{\mathbf{L}}_{r}$ as in Equation (17).
6:     Perform the minimization problem as in Equation (25).
7:     Compute learned generator matrix $\hat{\mathbf{L}}_{r}^{\alpha}$ by (27) and compare its properties to $\hat{\mathbf{L}}_{r}$.
8:Learning the full CG dynamics
9:     Learn effective potential $F^{\xi}$ by KDE or force matching (28).
10:     Effective drift is given by (24).
3.4 Overdamped Models for Molecular Systems
For the molecular examples considered in Sections 4.2 and 4.3, the dynamics are not governed by the reversible overdamped dynamics (3). Nevertheless, it can be expected that their position space dynamics are similar to the overdampled process after a re-scaling of time. This can be rigorously shown for the underdamped dynamics used in Section 4.2 37 . Fitting an overdamped model is simpler than modeling the full non-reversible dynamics if a non-linear CG map is used. Therefore, we compute the local diffusion as
$a^{\xi}_{\rm loc}(x)=\nabla_{x}\xi(x)\frac{2}{\beta\gamma}M^{-1}\nabla_{x}\xi^{\top}(x),$ (30)
where $M$ is the diagonal mass matrix of all atoms, $\beta$ is the inverse temperature, $\gamma$ is the friction, and $\nabla_{x}\xi(x)$ is the Jacobian of the CG map. This way of computing the local diffusion corresponds to a reversible overdamped Langevin process in full position space, meaning that the resulting CG model is an approximation to that same overdamped process. This also means that the eigenvalues and timescales of the overdamped model may differ from the ones obtained by directly approximating the full non-reversible process, but the meta-stable states will remain the same.
@@ 4 Examples
To show the effectiveness of the proposed method, we apply it to a two-dimensional model system defined by the Lemon-slice potential, and to MD simulation data of the alanine dipeptide and of the mini-protein Chignolin, which are widely-used test cases in molecular dynamics.
The optimization problems for learning the effective force and diffusion are solved by standard minimizers from scipy’s optimization toolbox. Note that in case the sought after object in the optimization blocks are decoupled, e.g. under the assumption that the effective diffusion is diagonal, we replace the minimizers with a simple least-square problem, making the step much cheaper.
@@ 4.1.1 System introduction
As shown in previous studies, the polar angle $\phi$ is a suitable CG coordinate for this system, as it resolves all four meta-stable states
@@
For this system, analytical expressions for the effective drift and diffusion along $\xi$ can be obtained by a slight modification of the results in Ref. 38 , and serve as reference values.
We apply our learning method with random Fourier features on the reaction coordinate $\xi$, to identify the generator eigenvalues and meta-stable states and, subsequently, to identify an effective dynamics along $\xi$ using Algorithm 1. As the polar angle is a periodic reaction coordinate (RC), we use the spectral measures associated to both a periodic and non-period Gaussian kernel and compare them. The kernel bandwidth in either versions of Gaussian kernel is optimized using cross validation based on the VAMP-score 26 . Details on the VAMP-score analysis are reported in the appendix.
@@ 4.1.2 Meta-stability analysis
Figure 1c shows the leading eigenvalues obtained from the generator matrix $\hat{\mathbf{L}}_{r}$. As one notices, there are four dominant eigenvalues followed by a gap. These four eigenvalues are corresponding to the four minima in the potential field. Having determined the eigenvectors of the generator, we can perform robust Perron Cluster Cluster Analysis (PCCA+) 39 algorithm to assign to each sample point its membership to each meta-stable state. Figure 1b shows that the four potential minima are perfectly recovered in this way.
@@
A comparison of the leading eigenvalues of the reference model $\hat{\mathbf{L}}_{r}$ and the learned matrix $\hat{\mathbf{L}}_{\alpha}^{\xi}$ for the optimal parameters $\alpha$ is shown in Figure 1c. Both choices of the kernel function lead to satisfactory results, the periodic kernel provides slightly higher accuracy in approximation of the generator eigenvalues. Note that the kernel bandwidth is tuned for each kernel function separately.
4.1.3 Analysis of the CG dynamics